# Why Machine Learning Models Systematically Underestimate Extreme Values II: How to Fix It with LatentNN

### v1

@@ Abstract
Attenuation bias—the systematic underestimation of regression coefficients due to measurement errors in input variables—affects astronomical data-driven models. For linear regression, this problem was solved by treating the true input values as latent variables to be estimated alongside model parameters. In this paper, we show that neural networks suffer from the same attenuation bias and that the latent variable solution generalizes directly to neural networks. We introduce LatentNN, a method that jointly optimizes network parameters and latent input values by maximizing the joint likelihood of observing both inputs and outputs. We demonstrate the correction on one-dimensional regression, multivariate inputs with correlated features, and stellar spectroscopy applications. LatentNN reduces attenuation bias across a range of signal-to-noise ratios where standard neural networks show large bias. This provides a framework for improved neural network inference in the low signal-to-noise regime characteristic of astronomical data. This bias correction is most effective when measurement errors are less than roughly half the intrinsic data range; in the regime of very low signal-to-noise and few informative features. Code is available at https://github.com/tingyuansen/LatentNN.
@@
I Introduction
Modern astronomy increasingly relies on neural networks for inference (Cranmer2020a; Huertas-Company2023; Ting2025). Data-driven approaches now routinely map observed quantities to physical parameters across a wide range of applications, from stellar parameters derived from spectra (e.g., Ness2015; Fabbro2018; Leung2019; Ting2019) to photometric redshifts from imaging (Hoyle2016; Li2022_GaZNet; Lin2022; Zhou2022) to stellar properties from time series light curves (Pasquet2019; Hon2021).
With the rise of neural network approaches, however, input uncertainties are often ignored or left unmodeled. This is largely a design choice rather than an oversight. Classical machine learning methods have long grappled with measurement errors in inputs, developing explicit frameworks for handling uncertain covariates. Neural network modeling, by contrast, arose in domains like computer vision (Simonyan2014; Szegedy2014; He2016) and natural language processing (Devlin2019; Brown2020; OpenAI2023; Touvron2023) where input noise is negligible, and its standard architectures have little mechanism for directly incorporating input uncertainties. When applied to astronomical data, this simplification can lead to systematic errors that are difficult to diagnose (Kelly2007; Ting2025_PaperI).
@@ I Introduction
This phenomenon is known as attenuation bias (Spearman1904; Frost2002; Kelly2007). In Paper I of this series (Ting2025_PaperI), we demonstrated this effect in detail for linear regression and explored its implications for astronomical data analysis. Attenuation bias manifests as a compression of the predicted dynamic range: high values are consistently predicted too low, while low values are predicted too high. The effect depends on the ratio of input measurement error to intrinsic signal range, and it is independent of sample size. As a rule of thumb, attenuation bias becomes important at input SNR of order 10, especially at low-dimensional problems. The exact magnitude depends on the dimensionality and correlation structure of the inputs, but the bias is never zero when input uncertainties are present. We refer interested readers to Paper I for detailed derivations and further discussion of attenuation bias in the context of linear regression.
@@
For linear models, the statistics literature has long recognized this problem and developed solutions such as Deming regression (Deming1943; Fuller1987; Carroll2006; Ting2025b). The key insight is to treat the true input values as latent variables to be estimated alongside model parameters (Loredo2004; Kelly2007). Rather than taking the noisy observations at face value, the model simultaneously infers what the true inputs likely were and learns the relationship between those true inputs and the outputs. The present paper extends this idea to neural networks. We show that neural networks suffer from the same attenuation bias as linear regression, and that the latent variable approach provides a general solution applicable to neural networks.
We introduce LatentNN, a practical implementation using modern automatic differentiation frameworks, and validate it on one-dimensional regression, multivariate inputs, and spectral applications. Section II briefly reviews attenuation bias following Paper I notation. Section III demonstrates that neural networks exhibit the same bias. Section IV presents the LatentNN formulation and results, first for the one-dimensional case, then extending to multivariate inputs. Section VI applies the method to stellar spectra. Section VII discusses limitations, extensions, and connections to hierarchical Bayesian modeling, followed by conclusions in Section VIII.
@@ II Attenuation Bias
This asymmetry between $x$ and $y$ arises because they play fundamentally different roles in regression: we predict $y$ given $x$, not the reverse. Errors in $x$ change the question being asked—we are predicting from a corrupted version of the input—which systematically weakens the apparent relationship. However, as we will see in Section IV, while $\sigma_{y}$ does not affect the bias itself, it does play a role in correcting the bias: the Deming regression solution (Deming1943; GolubVanLoan1980; Ting2025b) and its neural network generalization require specifying both $\sigma_{x}$ and $\sigma_{y}$ to properly weight the loss terms.
It is worth noting that this bias is sometimes mistakenly attributed to the last two causes. Observing that models underpredict extreme values, one might assume the problem lies in having too few training examples at the extremes or in noisy labels. However, attenuation bias has nothing to do with either of these issues. The bias arises purely from measurement errors in the input, and it persists even with perfectly balanced training data and perfect labels.
@@
III Attenuation Bias in Neural Networks
@@
This complexity makes it difficult to predict and therefore analytically correct the bias in closed form for general machine learning models. Here we first demonstrate empirically that multi-layer perceptrons (MLPs) exhibit the same qualitative behavior. MLPs, also known as fully connected neural networks (Cybenko1989; HornikStinchcombeWhite1989), are the simplest neural network architecture: each neuron in one layer is connected to every neuron in the next layer, with nonlinear activation functions between layers. Despite their simplicity, MLPs remain widely used in astronomy (e.g., Odewahn1992; Storrie-Lombardi1992; Collister2004; Guo2020; Wong2020; Wang2021; Li2022_LAMOST; Vynatheya2022; Winecki2024) and serve as the foundation for more complex architectures.
@@
For neural networks, we cannot directly measure a single coefficient $\beta$ as we did for linear regression in Section II. The MLP used here has over 4000 parameters distributed across multiple layers, rather than a single interpretable slope. Instead, we quantify attenuation through the predictions themselves. We define $\lambda_{y}$ as the slope of the regression of $y_{\rm pred}$ against $y_{\rm true}$. If a model suffers from attenuation bias, its predictions will be compressed toward the mean: when we plot $y_{\rm pred}$ versus $y_{\rm true}$, the slope will be less than unity. A value of $\lambda_{y}=1$ indicates no attenuation, while $\lambda_{y}<1$ indicates that predictions are systematically pulled toward the sample mean.
@@
Regularization does not resolve the issue either. One might hope that $L_{2}$ regularization (weight decay), which penalizes large weights, could counteract the attenuation (Tikhonov1963; HoerlKennard1970; Bishop2006; Ting2025b). However, regularization addresses overfitting by constraining model complexity, not systematic bias in the inputs. The bias arises from the mismatch between $\mathbf{x}_{\rm obs}$ and $\mathbf{x}_{\rm true}$, which regularization cannot correct. We verified this empirically by sweeping over a range of $L_{2}$ regularization strengths: the attenuation factor $\lambda_{y}$ remains unchanged regardless of the regularization parameter.
IV LatentNN: Correcting Attenuation Bias with Latent Variables
Having established that neural networks suffer from attenuation bias, we now develop a method to correct it. The approach builds on a classical idea from the statistics literature on errors-in-variables regression (Carroll1995; Carroll2006; Bishop2006): rather than treating noisy observations as exact, we treat the unknown true values as latent variables to be estimated alongside the model parameters. For linear regression, this leads to Deming regression (Deming1943). Here we show that the same principle extends naturally to neural networks, yielding a practical correction method we call LatentNN.
IV.1 The Latent Variable Solution in Linear Regression
@@
This approach of treating unknown true values as parameters to be inferred is an example of hierarchical modeling (Loredo2004; Kelly2007; Gelman2013). In this hierarchical structure, the model parameter $\beta$ sits at the top level, the true values $x_{{\rm true},i}$ form the intermediate level, and the actual measurements $(x_{\rm obs},y_{\rm obs})$ sit at the bottom level. The model constrains the intermediate level, while the intermediate level generates the observations.
@@
where $S_{xx}=\sum_{i}(x_{{\rm obs},i}-\bar{x})^{2}$, $S_{yy}=\sum_{i}(y_{{\rm obs},i}-\bar{y})^{2}$, and $S_{xy}=\sum_{i}(x_{{\rm obs},i}-\bar{x})(y_{{\rm obs},i}-\bar{y})$ are the sum of squared deviations and cross-products, with $\bar{x}$ and $\bar{y}$ denoting the sample means. In the limit $\sigma_{x}\to 0$ (equivalently $r\to\infty$), this reduces to the ordinary least squares estimator $\hat{\beta}_{\rm OLS}=S_{xy}/S_{xx}$. When $\sigma_{x}>0$, the Deming estimator yields a steeper slope than OLS, correcting for the attenuation (see also Ting2025_PaperI).
@@ IV.2 The LatentNN Formulation
$\mathcal{L}(\bm{\theta},\{x_{{\rm latent},i}\})$ $=\sum_{i=1}^{N}\frac{(y_{{\rm obs},i}-f_{\bm{\theta}}(x_{{\rm latent},i}))^{2}}{\sigma_{y,i}^{2}}$
@@
$\hat{\bm{\theta}},\{\hat{x}_{{\rm latent},i}\}=\underset{\bm{\theta},\{x_{{\rm latent},i}\}}{\arg\min}\;\mathcal{L}(\bm{\theta},\{x_{{\rm latent},i}\}).$ (12)
@@ IV.3 Implementation
$\Phi^{(t+1)}=\Phi^{(t)}-\eta\nabla_{\Phi}\mathcal{L},$ (13)
where $\eta$ is the learning rate and $\nabla_{\Phi}\mathcal{L}$ is the gradient of the loss with respect to all parameters. This is the same update rule used in standard neural network training, but with $\Phi$ now including both the network weights $\bm{\theta}$ and the latent inputs $\{x_{{\rm latent},i}\}$. This represents a conceptual shift from typical neural network applications, where we optimize only over $\bm{\theta}$. Here, we treat the observed inputs as noisy measurements rather than ground truth, and optimize jointly over $\bm{\theta}$ and the latent inputs.
In modern deep learning frameworks such as PyTorch or JAX, this is straightforward to implement. The latent values $x_{{\rm latent},i}$ are declared as additional learnable parameters, and the framework automatically computes gradients via backpropagation (RumelhartHintonWilliams1986; Lecun1998).
Several practical considerations affect performance. First, initialization matters. We initialize the latent values at the observed values, $x_{{\rm latent},i}^{(0)}=x_{{\rm obs},i}$, which ensures that the initial loss is reasonable and provides a starting point consistent with our probabilistic model.
@@ V Numerical Experiments
V.1 One-Dimensional Case
We use the same data setup and network architecture as in Section III: the true relationship $y=2x_{\rm true}$ with 1000 training samples, and an MLP with 2 hidden layers of 64 units trained for 20000 epochs. For LatentNN, we add a learnable parameter $x_{{\rm latent},i}$ for each training sample, initialized to $x_{{\rm obs},i}$.
@@ V.1 One-Dimensional Case
Figure 3 shows a representative comparison at $\mathrm{SNR}_{x}=1$, using $\mathrm{SNR}_{y}=10$ as in Section III. At $\mathrm{SNR}_{x}=1$, the measurement error equals the signal range ($\sigma_{x}=\sigma_{\rm range}$)—an extreme regime where the noise is 100% of the data’s dynamic range. The top panels show predicted versus true values on the held-out test set. The standard MLP shows $\lambda_{y}\approx 0.5$, demonstrating large attenuation bias. LatentNN recovers $\lambda_{y}\approx 1$, correcting the bias. The bottom-left panel shows the learned functions: the standard MLP learns an attenuated slope, while LatentNN recovers the true relationship $f(x)=2x$.
@@
To systematically evaluate performance, we repeat each experiment 8 times with different random seeds, using the same three-fold data split (train/validation/test) for each run. Figure 4 shows $\lambda_{y}$ versus $\mathrm{SNR}_{x}$ across a range of noise levels from $\mathrm{SNR}_{x}=1$ to $\mathrm{SNR}_{x}=10$, testing three $\mathrm{SNR}_{y}$ values (3, 10, and 30). As expected from the theory in Section II, varying $\mathrm{SNR}_{y}$ has little effect on attenuation—the bias depends primarily on input noise. The standard MLP follows the theoretical attenuation curve $\lambda_{y}=1/(1+(\sigma_{x}/\sigma_{\rm range})^{2})$ regardless of $\mathrm{SNR}_{y}$. LatentNN maintains $\lambda_{y}\approx 1$ across all tested $\mathrm{SNR}_{x}$ and $\mathrm{SNR}_{y}$ values. LatentNN provides robust correction even at $\mathrm{SNR}_{x}=1$, though with increased run-to-run scatter at lower $\mathrm{SNR}_{x}$ values as the problem becomes increasingly ill-conditioned.
@@
V.2 Multivariate Inputs
Real astronomical applications involve multivariate inputs. The LatentNN formulation generalizes directly: the latent variables $\mathbf{X}_{\rm latent}$ form an $N\times p$ matrix, where $p$ is the number of input features. The loss function from Equation 11 extends naturally to
$\mathcal{L}$ $=\sum_{i=1}^{N}\frac{(y_{i}-f_{\bm{\theta}}(\mathbf{X}_{{\rm latent},i}))^{2}}{\sigma_{y,i}^{2}}$

### v2

@@ Abstract
Attenuation bias—the systematic underestimation of regression coefficients due to measurement errors in input variables—affects astronomical data-driven models. For linear regression, this problem was solved by treating the true input values as latent variables to be estimated alongside model parameters. In this paper, we show that neural networks suffer from the same attenuation bias and that the latent variable solution generalizes directly to neural networks. We introduce LatentNN, a method that jointly optimizes network parameters and latent input values by maximizing the joint likelihood of observing both inputs and outputs. We demonstrate the correction on one-dimensional regression, multivariate inputs with correlated features, and stellar spectroscopy applications. LatentNN reduces attenuation bias across a range of signal-to-noise ratios where standard neural networks show large bias. This provides a framework for improved neural network inference in the low signal-to-noise regime characteristic of astronomical data. This bias correction is most effective when measurement errors are less than roughly half the intrinsic data range, in the regime of very low signal-to-noise and few informative features. Code is available at https://github.com/tingyuansen/LatentNN.
keywords:
@@
1 Introduction
Modern astronomy increasingly relies on neural networks for inference (Cranmer et al., 2020; Huertas-Company and Lanusse, 2023; Ting, 2025a). Data-driven approaches now routinely map observed quantities to physical parameters across a wide range of applications, from stellar parameters derived from spectra (e.g., Ness et al., 2015; Fabbro et al., 2018; Leung and Bovy, 2019; Ting et al., 2019) to photometric redshifts from imaging (Hoyle, 2016; Li et al., 2022a; Lin et al., 2022; Zhou et al., 2022) to stellar properties from time series light curves (Pasquet et al., 2019; Hon et al., 2021).
With the rise of neural network approaches, however, input uncertainties are often ignored or left unmodeled. This is largely a design choice rather than an oversight. Classical machine learning methods have long grappled with measurement errors in inputs, developing explicit frameworks for handling uncertain covariates. Neural network modeling, by contrast, arose in domains like computer vision (Simonyan and Zisserman, 2014; Szegedy et al., 2014; He et al., 2016) and natural language processing (Devlin et al., 2018; Brown et al., 2020; OpenAI et al., 2023; Touvron et al., 2023) where input noise is negligible, and its standard architectures have little mechanism for directly incorporating input uncertainties. When applied to astronomical data, this simplification can lead to systematic errors that are difficult to diagnose (Kelly, 2007; Ting, 2025c).
@@ 1 Introduction
This phenomenon is known as attenuation bias (Spearman, 1904; Frost and Thompson, 2002; Kelly, 2007). In Paper I of this series (Ting, 2025c), we demonstrated this effect in detail for linear regression and explored its implications for astronomical data analysis. Attenuation bias manifests as a compression of the predicted dynamic range: high values are consistently predicted too low, while low values are predicted too high. The effect depends on the ratio of input measurement error to intrinsic signal range, and it is independent of sample size. As a rule of thumb, attenuation bias becomes important at input SNR of order 10, especially at low-dimensional problems. The exact magnitude depends on the dimensionality and correlation structure of the inputs, but the bias is never zero when input uncertainties are present. We refer interested readers to Paper I for detailed derivations and further discussion of attenuation bias in the context of linear regression.
@@
For linear models, the statistics literature has long recognized this problem and developed solutions such as Deming regression (Deming, 1943; Fuller, 1987; Carroll et al., 2006; Ting, 2025b). The key insight is to treat the true input values as latent variables to be estimated alongside model parameters (Loredo, 2004; Kelly, 2007). Rather than taking the noisy observations at face value, the model simultaneously infers what the true inputs likely were and learns the relationship between those true inputs and the outputs. The present paper extends this idea to neural networks. We show that neural networks suffer from the same attenuation bias as linear regression, and that the latent variable approach provides a general solution applicable to neural networks.
We introduce LatentNN, a practical implementation using modern automatic differentiation frameworks, and validate it on one-dimensional regression, multivariate inputs, and spectral applications. Section 2 briefly reviews attenuation bias following Paper I notation. Section 3 demonstrates that neural networks exhibit the same bias. Section 4 presents the LatentNN formulation and results, first for the one-dimensional case, then extending to multivariate inputs. Section 6 applies the method to stellar spectra. Section 7 discusses limitations, extensions, and connections to hierarchical Bayesian modeling, followed by conclusions in Section 8.
@@ 2 Attenuation Bias
This asymmetry between $x$ and $y$ arises because they play fundamentally different roles in regression: we predict $y$ given $x$, not the reverse. Errors in $x$ change the question being asked—we are predicting from a corrupted version of the input—which systematically weakens the apparent relationship. However, as we will see in Section 4, while $\sigma_{y}$ does not affect the bias itself, it does play a role in correcting the bias: the Deming regression solution (Deming, 1943; Golub and Van Loan, 1980; Ting, 2025b) and its neural network generalization require specifying both $\sigma_{x}$ and $\sigma_{y}$ to properly weight the loss terms.
It is worth noting that this bias is sometimes mistakenly attributed to insufficient training data at the extremes or to noisy labels. Observing that models underpredict extreme values, one might assume the problem lies in having too few training examples at the extremes or in noisy labels. However, attenuation bias has nothing to do with either of these issues. The bias arises purely from measurement errors in the input, and it persists even with perfectly balanced training data and perfect labels.
@@
3 Attenuation Bias in Neural Networks
@@
This complexity makes it difficult to predict and therefore analytically correct the bias in closed form for general machine learning models. Here we first demonstrate empirically that multi-layer perceptrons (MLPs) exhibit the same qualitative behavior. MLPs, also known as fully connected neural networks (Cybenko, 1989; Hornik et al., 1989), are the simplest neural network architecture: each neuron in one layer is connected to every neuron in the next layer, with nonlinear activation functions between layers. Despite their simplicity, MLPs remain widely used in astronomy (e.g., Odewahn et al., 1992; Storrie-Lombardi et al., 1992; Collister and Lahav, 2004; Guo et al., 2020; Wong et al., 2020; Wang et al., 2021; Li et al., 2022b; Vynatheya et al., 2022; Winecki and Kochanek, 2024) and serve as the foundation for more complex architectures.
@@
For neural networks, we cannot directly measure a single coefficient $\beta$ as we did for linear regression in Section 2. The MLP used here has over 4000 parameters distributed across multiple layers, rather than a single interpretable slope. Instead, we quantify attenuation through the predictions themselves. Here, predictions are always made using the observed (noisy) inputs: $y_{\rm pred}=f_{\bm{\theta}}(x_{\rm obs})$, not the unknown true inputs. We define $\lambda_{y}$ as the slope of the regression of $y_{\rm pred}$ against $y_{\rm true}$. If a model suffers from attenuation bias, its predictions will be compressed toward the mean: when we plot $y_{\rm pred}$ versus $y_{\rm true}$, the slope will be less than unity. A value of $\lambda_{y}=1$ indicates no attenuation, while $\lambda_{y}<1$ indicates that predictions are systematically pulled toward the sample mean.
@@
Regularization does not resolve the issue either. One might hope that $L_{2}$ regularization (weight decay), which penalizes large weights, could counteract the attenuation (Tikhonov, 1963; Hoerl and Kennard, 1970; Bishop, 2006; Ting, 2025b). However, regularization addresses overfitting by constraining model complexity, not systematic bias in the inputs. The bias arises from the mismatch between $\mathbf{x}_{\rm obs}$ and $\mathbf{x}_{\rm true}$, which regularization cannot correct. We verified this empirically by sweeping over a range of $L_{2}$ regularization strengths: the attenuation factor $\lambda_{y}$ remains unchanged regardless of the regularization parameter.
4 LatentNN: Correcting Attenuation Bias with Latent Variables
Having established that neural networks suffer from attenuation bias, we now develop a method to correct it. The approach builds on a classical idea from the statistics literature on errors-in-variables regression (Carroll et al., 1995, 2006; Bishop, 2006): rather than treating noisy observations as exact, we treat the unknown true values as latent variables to be estimated alongside the model parameters. For linear regression, this leads to Deming regression (Deming, 1943). Here we show that the same principle extends naturally to neural networks, yielding a practical correction method we call LatentNN.
4.1 The Latent Variable Solution in Linear Regression
@@
Here $\sigma_{y}^{2}$ is the known measurement uncertainty in the output. In principle, one could add a jitter term to account for intrinsic scatter about the true relationship, but we assume pure measurement noise for simplicity.
@@
This approach of treating unknown true values as parameters to be inferred is an example of hierarchical modeling (Loredo, 2004; Kelly, 2007; Gelman et al., 2013). In this hierarchical structure, the model parameter $\beta$ sits at the top level, the true values $x_{{\rm true},i}$ form the intermediate level, and the actual measurements $(x_{\rm obs},y_{\rm obs})$ sit at the bottom level. The model constrains the intermediate level, while the intermediate level generates the observations.
@@
where $S_{xx}=\sum_{i}(x_{{\rm obs},i}-\bar{x})^{2}$, $S_{yy}=\sum_{i}(y_{{\rm obs},i}-\bar{y})^{2}$, and $S_{xy}=\sum_{i}(x_{{\rm obs},i}-\bar{x})(y_{{\rm obs},i}-\bar{y})$ are the sum of squared deviations and cross-products, with $\bar{x}$ and $\bar{y}$ denoting the sample means. In the limit $\sigma_{x}\to 0$ (equivalently $r\to\infty$), this reduces to the ordinary least squares estimator $\hat{\beta}_{\rm OLS}=S_{xy}/S_{xx}$. When $\sigma_{x}>0$, the Deming estimator yields a steeper slope than OLS, correcting for the attenuation (see also Ting, 2025c).
@@ 4.2 The LatentNN Formulation
$\mathcal{J}(\bm{\theta},\{x_{{\rm latent},i}\})$ $=\sum_{i=1}^{N}\frac{(y_{{\rm obs},i}-f_{\bm{\theta}}(x_{{\rm latent},i}))^{2}}{\sigma_{y,i}^{2}}$
@@
$\hat{\bm{\theta}},\{\hat{x}_{{\rm latent},i}\}=\underset{\bm{\theta},\{x_{{\rm latent},i}\}}{\arg\min}\;\mathcal{J}(\bm{\theta},\{x_{{\rm latent},i}\}).$ (12)
@@ 4.3 Implementation
$\Phi^{(t+1)}=\Phi^{(t)}-\eta\nabla_{\Phi}\mathcal{J},$ (13)
where $\eta$ is the learning rate and $\nabla_{\Phi}\mathcal{J}$ is the gradient of the loss with respect to all parameters. This is the same update rule used in standard neural network training, but with $\Phi$ now including both the network weights $\bm{\theta}$ and the latent inputs $\{x_{{\rm latent},i}\}$. This represents a conceptual shift from typical neural network applications, where we optimize only over $\bm{\theta}$. Here, we treat the observed inputs as noisy measurements rather than ground truth, and optimize jointly over $\bm{\theta}$ and the latent inputs.
In modern deep learning frameworks such as PyTorch or JAX, this is straightforward to implement. The latent values $x_{{\rm latent},i}$ are declared as additional learnable parameters, and the framework automatically computes gradients via backpropagation (Rumelhart et al., 1986; LeCun et al., 1998).
Several practical considerations affect performance. First, initialization matters. We initialize the latent values at the observed values:
$x_{{\rm latent},i}^{(0)}=x_{{\rm obs},i}.$ (14)
In practice, this means declaring a learnable parameter array of shape $(N,p)$ initialized to the observed input matrix. This ensures that the initial loss is reasonable and provides a starting point consistent with our probabilistic model.
@@ 5 Numerical Experiments
5.1 One-Dimensional Case
We use the same data setup and network architecture as in Section 3: the true relationship $y=2x_{\rm true}$ with 1000 training samples, and an MLP with 2 hidden layers of 64 units trained for 20000 epochs. For LatentNN, we add a learnable parameter $x_{{\rm latent},i}$ for each training sample, initialized to $x_{{\rm obs},i}$.
@@ 5.1 One-Dimensional Case
Figure 3 shows a representative comparison at $\mathrm{SNR}_{x}=1$, using $\mathrm{SNR}_{y}=10$ as in Section 3. At $\mathrm{SNR}_{x}=1$, the measurement error equals the signal range ($\sigma_{x}=\sigma_{\rm range}$)—an extreme regime where the noise is 100% of the data’s dynamic range. The top panels show predicted versus true values on the held-out test set. The standard MLP shows $\lambda_{y}\approx 0.5$, demonstrating large attenuation bias. LatentNN recovers $\lambda_{y}\approx 1$, correcting the bias. The bottom-left panel shows the learned functions: the standard MLP learns an attenuated slope, while LatentNN recovers the true relationship $f(x)=2x$.
@@
To systematically evaluate performance, we repeat each experiment 8 times with different random seeds, using the same three-fold data split (train/validation/test) for each run. Figure 4 shows $\lambda_{y}$ versus $\mathrm{SNR}_{x}$ across a range of noise levels from $\mathrm{SNR}_{x}=1$ to $\mathrm{SNR}_{x}=10$, testing three $\mathrm{SNR}_{y}$ values (3, 10, and 30). As expected from the theory in Section 2, varying $\mathrm{SNR}_{y}$ has little effect on attenuation—the bias depends primarily on input noise. The standard MLP follows the theoretical attenuation curve $\lambda_{y}=1/(1+(\sigma_{x}/\sigma_{\rm range})^{2})$ regardless of $\mathrm{SNR}_{y}$. LatentNN maintains $\lambda_{y}\approx 1$ across all tested $\mathrm{SNR}_{x}$ and $\mathrm{SNR}_{y}$ values. LatentNN provides robust correction even at $\mathrm{SNR}_{x}=1$, though with increased run-to-run scatter at lower $\mathrm{SNR}_{x}$ values as the problem becomes increasingly ill-conditioned.
@@
5.2 Multivariate Inputs
Real astronomical applications involve multivariate inputs—for example, stellar spectra contain many flux measurements that serve as inputs to predict stellar parameters. The LatentNN formulation generalizes directly: the latent variables $\mathbf{X}_{\rm latent}$ form an $N\times p$ matrix, where $p$ is the number of input features. The loss function from Equation 11 extends naturally to
$\mathcal{J}$ $=\sum_{i=1}^{N}\frac{(y_{i}-f_{\bm{\theta}}(\mathbf{X}_{{\rm latent},i}))^{2}}{\sigma_{y,i}^{2}}$